### v2

# Style Ambiguity Loss Without Labeled Datasetss

Abstract
In this work, we explore using the style ambiguity training objective, originally used to approximate creativity, on a diffusion model. However, this objective requires the use of a pretrained classifier and a labeled dataset. We introduce new forms of style ambiguity loss that do not require training a classifier or a labeled dataset, and show that our new methods score higher both on automated metrics and user studies to analyze novelty and appreciation. Code available at https://github.com/jamesBaker361/clipcreate
1 Introduction
Much of generative imagery is trained to mimic the underlying dataset. However, there is also demand for machine-generated images that are not just realistic but artistic, as exemplified by the picture that sold for nearly half a million dollars (Christie’s, 2018) at auction, or the picture used in the sitcom Silicon Valley (AICAN, 2024). While the definition of art is somewhat philosophical and beyond the scope of this paper, it most certainly has to be creative. Creative assets are usually defined as being "novel and useful" (Diedrich et al., 2015). A string of random characters is novel in that we cannot predict the next characters. However, if the purpose of characters is to compose words that compose sentences that communicate a message, then a string of random characters is not useful. Most people would not consider a string of random characters to be creative or art. If the utility of an image is to depict objects and scenes that humans understand and can recognize, then generative models perform well in that regard. Novelty, on the other hand has been underexplored. A breakthrough was the invention of the Creative Adversarial Network (Elgammal et al., 2017), which used a style ambiguity loss to train a generator network to generate images that could not be classified as belonging to a particular style. However, this style ambiguity loss requires a pretrained classifier: Every set of styles or concepts requires training a classifier before even training a model to generate images, which itself takes time and requires a labeled dataset, the collection and curation of which can be expensive. To circumvent these issues, we propose using a classifier for style ambiguity that does not require any additional training and can be easily applied to any dataset, labeled or unlabeled. We demonstrate this by using reinforcement learning to finetune diffusion models with the various style ambiguity losses. Our contributions are as follows:
  • We developed versatile CLIP-based and K-Means-based creative style ambiguity losses that do not require training a separate style classifier on a labeled dataset.
  • We find that diffusion models trained with our novel forms of style ambiguity loss score similar or higher on automated metrics of human appreciation to diffusion models trained without, and perform a user study to show that our novel models also score higher on components of creativity.
2 Related Work
2.1 Creativity
Creative work has been formulated as work having novelty, in that it differs from other similar objects, and also utility, in that it still performs a function (Cropley, 2006). For example, a Corinthian column has elaborate, interesting, unexpected adornments (novelty) but still holds up a building (utility). A distinction can also be made between "P-creativity", where the work is novel to the creator, and "H-creativity" where the work is novel to everyone (Boden, 1990), the latter being of interest. There have been many attempts to make models that can be "H-creative". Such techniques include using genetic algorithms (DiPaola & Gabora, 2008), reconstructing artifacts from novel collections of attributes (Iqbal et al., 2016), and most relevantly to this work, using Generative Adversarial Networks (Elgammal et al., 2017; Nobari et al., 2021).
2.2 Computational Art
One of the first algorithmic approaches dates back to the 1970s with the now primitive AARON (McCorduck, 1991), which was initially only capable of drawing black and white sketches. Generative Adversarial Networks (Goodfellow et al., 2014), or GANs, were some of the first models to be able to create complex, photorealistic images and seemed to have potential to be able to make art. Despite many problems with GANs, such as mode collapse and unstable training (Saxena & Cao, 2023), GANs and further improvements (Arjovsky et al., 2017; Karras et al., 2019; 2018) were state of the art until the introduction of diffusion Sohl-Dickstein et al. (2015). These models, such as IMAGEN (Saharia et al., 2022), DALLE-3 (Betker et al., ) and the Stable Diffusion family (Rombach et al., 2022b; Esser et al., 2024; Podell et al., 2023) have attained widespread adoption due to their ability to generate high-quality images from text prompts.
2.3 Reinforcement Learning
Reinforcement learning (RL) is a method of training a model by having it take actions that generate a reward signal and change the environment, thus changing the impact and availability of future actions Qiang & Zhongli (2011). RL has been used for tasks as diverse as playing board games (Silver et al., 2017), protein design (Lutz et al., 2023), self-driving vehicles (Kiran et al., 2021) and quantitative finance (Sahu et al., 2023). Policy-gradient RL (Sutton et al., 1999) optimizes a policy $\pi$ that chooses which action to take at any given timestep, as opposed to value-based methods that may use a heuristic to determine the optimal choice. Examples of policy gradient methods include Soft Actor Critic (Haarnoja et al., 2018), Deep Deterministic Policy Gradient (Lillicrap et al., 2019) and Trust Region Policy Optimization (Schulman et al., 2017a).
3 Background
3.1 Creative Adversarial Network
A Generative Adversarial Network, or GAN (Goodfellow et al., 2014), consists of two models, a generator and a discriminator. The generator generates samples from noise, and the discriminator detects if the samples are drawn from the real data or generated. During training, the generator is trained to trick the discriminator into classifying generated images as real, and the discriminator is trained to classify images correctly. Given a generator $G:\mathbb{R}^{noise}\rightarrow\mathbb{R}^{h\times w\times 3}$, a discriminator $D:\mathbb{R}^{h\times w\times 3}\rightarrow[0,1]$ real images $x\in\mathbb{R}^{h\times w\times 3}$, and noise $\mathcal{Z}\in\mathbb{R}^{noise}$, the objective is:
[…]
Elgammal et al. (2017) introduced the Creative Adversarial Network, or CAN, which was a DCGAN (Radford et al., 2016) where the discriminator was also trained to classify real samples, minimizing the style classification loss. Given $N$ classes of image (such as ukiyo-e, baroque, impressionism, etc.), the classification modules of the Discriminator $D_{C}:\mathbb{R}^{h\times w\times 3}\rightarrow\mathbb{R}^{N}$ that returns a probability distribution over the $N_{s}$ style classes for an image and the real labels $\ell\in\mathbb{R}^{N}$, the style classification loss was:
$L_{SL}=\mathbb{E}_{x,\ell}[\textbf{CE}(D_{C}(x),\ell)]$
Where CE is the cross entropy function.
The generator was also trained to generate samples that could not be easily classified as belonging to one class. This stylistic ambiguity is a proxy for creativity or novelty. Given a vector $U\in\mathbb{R}^{N}$, where each entry $u_{1},u_{2},,,u_{N}=\frac{1}{N}$, and the classification modules of the discriminator $D_{C}$ the style ambiguity loss is:
$L_{SA}=\mathbb{E}_{\mathcal{Z}}[\textbf{CE}(C(G(\mathcal{Z})),U)]$
The discriminator was additionally trained to minimize $L_{SL}$ and the generator was additionally trained to minimize $L_{SA}$.
3.2 Diffusion
A diffusion model aims to learn to iteratively remove the noise from a corrupted sample to restore the original. Starting with $x_{0}$, the forward process $q$ iteratively adds Gaussian noise to produce the noised version $x_{T}$, using a noise schedule $\beta_{1}...\beta_{T}$, which can be learned or manually set as a hyperparameter:
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I})$
More importantly, we also want to model the reverse process $p$, that turns a noisy sample $x_{T}$ back into $x_{0}$, conditioned on some context $c$. As $x_{T}$ is the fully noised version, $p(x_{T}|c)=\mathcal{N}(x_{T};\mathbf{0},\mathbf{I})$
$p_{\theta}(x_{t-1}|x_{t},c)=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t,c),\Sigma_{\theta}(x_{t},t,c))$
Once the model has been trained, the reverse process, aka inference, generates a sample from noise $x_{T}\sim\mathcal{N}(0,1)$. We use the DDIM technique (Song et al., 2022) for sampling. In this work, we the more popular variant, Stable Diffusion (Rombach et al., 2022a), where $x$ is replaced with a latent embedding $\mathcal{E}(x)$, where $\mathcal{E}:\mathbb{R}^{h\times w\times 3}\rightarrow\mathbb{R}^{h_{z}\times w_{z}\times c_{z}}$ is a frozen autoencoder (Kingma & Welling, 2022), and $h_{z}<h,w_{z}<w,c_{z}>3$.
4 Methods
4.1 Denoising Diffusion Proximal Optimisation
Introduced by Black et al. (2023), Denoising Diffusion Proximal Optimisation, or DDPO, represents the reverse Diffusion Process as a Markov Decision Process (Bellman, 1957). A similar method was also pursued by Fan et al. (2023). Reinforcement learning training was then applied to a pretrained diffusion model. Following Schulman et al. (2017b), Black et al. (2023) also implemented clipping to protect the policy gradient from excessively large updates, and per prompt stat tracking to normalize rewards. We largely follow their method but use a different reward function. We fine-tune off of the pre-existing stabilityai/stable-diffusion-2-base checkpoint (Rombach et al., 2022a) downloaded from https://huggingface.co/stabilityai/stable-diffusion-2-base.
4.2 Text Prompts
DDPO does not require any new data, given that we are fine-tuning off of a pretrained checkpoint. However, each time we train the model, we must decide which text prompts to use to condition the generation of images. We used the set of (painting, drawing, art) as our text prompts.
4.3 Datasets and Labels
Training the CAN, of course, requires a labeled dataset. We used two different real datasets based off of WikiArt (Saleh & Elgammal, 2015):
  1. 1.
    Full: the WikiArt dataset as is. Consists of roughly 80k images.
  2. 2.
    Mediums: We used BLIP (Li et al., 2022) to generate captions, and then selected the 10 classes that had the highest fraction of their descriptions containing any of the words in the text prompt set (painting, drawing, art). We then used the images in WikiArt that belonged to those classes. Consists of roughly 20k images.
This also meant we had two different sets of style class labels: $L_{full}$, all 27 class labels, used with Full and $L_{med}$, the 10 labels used in Mediums. For implementation details regarding the captions, and a list of the class labels used for Mediums refer to appendix D
4.4 DDPO Reward Function
In the original DDPO paper, the authors used four different reward functions for four different tasks. For example, they used a scorer trained on the LAION dataset (Schuhmann & Beaumont, 2022) as the reward function to improve the aesthetic quality of generated outputs. In this paper, we use the reward model based on Elgammal et al. (2017), where the model is rewarded for stylistic ambiguity, combined with a reward for utility. Given a pretrained CLIP (Radford et al., 2021) model, that can return a similarity score for each image-text pair: $CLIP:\mathbb{R}^{text}\times\mathbb{R}^{h\times w\times 3}\rightarrow\mathbb{R}$, image $x_{0}\in\mathbb{R}^{h\times w\times 3}$ generated with text prompt $s$, cross entropy CE, uniform distribution $U\in\mathbb{R}^{N}$ and a classifier $C:\mathbb{R}^{h\times w\times 3}\rightarrow\mathbb{R}^{N}$ we want to maximize:
$R(x_{0})=-\lambda_{novelty}\textbf{CE}(C(x_{0}),U)+\lambda_{utility}CLIP(s,x_{0})$
The first term on the left side of the equation represents style ambiguity loss, and the second term maintains alignment between text and image, essentially keeping the model from straying "too far" from the text prompt; these terms approximate novelty and utility, respectively. We actually have multiple choices of classifier, which we discuss. The Discriminator Classifier is what was used for the CAN, and the CLIP-Based Style Classifier and K-Means Image Based Classifer have never been used for style ambiguity loss before, and our contribution is to introduce them and show their merits.
4.4.1 Discriminator Classifier
We can use the classification module of the CAN discriminator as the classifier in the reward function, setting $C=D_{C}$. This is the traditional method of style ambiguity loss, and the baseline against which we are comparing the other two types of classifier with. This discriminator was trained at image resolution 512, with batch size 128 without using gradient penalty (Gulrajani et al., 2017). Given that we had two datasets (Full and Mediums), we actually had to train two of these discriminators, one trained for each. The disadvantage of using a discriminator is that it requires a labeled dataset, which are expensive and time-consuming to create. Even the use of VLMs (Ghosh et al., 2024) for labeling may not suffice, given that expert domain knowledge is required in many cases. We thus introduce two methods of style ambiguity that do not require a labeled dataset.
4.4.2 CLIP-Based Style Classifier
For each generated image $x_{0}$, for each class name $s_{i},1\leq i\leq N_{s}$ in the style class label set we want to use, we find $CLIP(s_{i},x_{0})$. We can then create a vector $(CLIP(s_{1},x_{0}),CLIP(s_{2},x_{0}),,,CLIP(s_{N_{s}},x_{0}))$ and then use softmax to normalize the vector and define the result as $CC(x_{0})$. Formally:
$CC(x_{0})=\textbf{softmax}((CLIP(s_{1},x_{0}),CLIP(s_{2},x_{0}),,,CLIP(s_{N_{s}},x_{0}))$
Then we set $C=CC$. Given the two sets of labels, $L_{full}$ and $L_{med}$, we had two different types of CLIP-Based classifier.
4.4.3 K-Means Image Based Classifiers
Alternatively, we can embed the images (In our case, we embedded them into the CLIP embedding space $\in\mathbb{R}^{768}$) and perform k-means clustering to generate k centers. Given a CLIP Embedder $E:\mathbb{R}^{h\times w\times 3}\rightarrow\mathbb{R}^{768}$ mapping images to embeddings, and the k centers $c_{1},c_{2},,,c_{k}$ we can create a vector $(\frac{1}{||E(x_{0})-c_{1}||},\frac{1}{||E(x_{0})-c_{2}||},,,,\frac{1}{||E(x_{0})-c_{k}||}$ and then use softmax to normalize the vector and define the result as $KM(x_{0}))$. Formally:
$KMEANS(x_{0})=\textbf{softmax}(\frac{1}{||E(x_{0})-c_{1}||},\frac{1}{||E(x_{0})-c_{2}||},,,,\frac{1}{||E(x_{0})-c_{k}||})$
Then we set $C=KM$. Given two datasets (Full and Mediums), we had two sets of centers.
5 Results
Given three types of classifier and two datasets, we had six models.
We generated all images with width and height = 512. The authors used width and height = 256 in the original CAN paper. However, given that larger, more detailed images are preferred by most people, we thought it more relevant to focus on larger images.
5.1 Automated Evaluation
For each dataset, for each choice of classifier, we made an evaluation dataset of 200 images. We used the exact same prompts and random seeds for each. I.e. if the nth image generated by disc-full used prompt "painting" and random seed = $z\in\mathbb{Z}$, so would the nth image generated by clip-full, kmeans-full, etc. for prompts, we had 24 prompts, each generated by concatenating one of ["painting","art","drawing"] with one of [" of a man ", " of a woman "," of nature "," of an animal ", " of a city"," "," of a building "," of a child "] We used two scoring metrics to score the models. Both of these metrics are metrics of human appreciation:
  • AVA Score: (AVA) The AVA model was trained on the AVA dataset (Murray et al., 2016) of images and average rankings by human subjects, in order to learn to approximate human preferences given an image. We used the CLIP model weights from the clip-vit-large-patch14 checkpoint and the Multi-Layer Perceptron weights downloaded from https://huggingface.co/trl-lib/ddpo-aesthetic-predictor.
  • Image Reward: (IR) The image reward model (Xu et al., 2023) was trained to score images given their text description based on a dataset of images and human rankings. We used the image-reward python library found at https://github.com/THUDM/ImageReward/tree/main.
Model AVA IR
$disc-full$ 4.96 (0.5) -1.26 (1.04)
$clip-full$ 4.99 (0.46) -1.59 (0.63)
$kmeans-full$ 5.71 (0.42) 0.03 (1.35)
$disc-med$ 5.12 (0.4) -1.55 (0.59)
$clip-med$ 5.44 (0.41) 0.91 (0.89)
$kmeans-med$ 5.09 (0.67) 0.77 (0.87)
Table 1: Aesthetic Scores, Diffusion Models
[Uncaptioned image] [Uncaptioned image]
Table 2: Score Box Plots
Results of our experiments are shown in Table 1 Each cell contains the mean (and standard deviation). Box plots are shown in 2. We can see that there was less variance and generally less skew for the aesthetic score than for the Image Reward. For both datasets, for both metrics, either a clip model or kmeans model scored higher than the baseline discriminator model. This shows that our discriminator-free approaches in do not come at the cost of lower quality, and in fact are likely to produce better images than the discriminator baseline.
5.1.1 Comparison with GAN Baseline
GANs are largely out of fashion, but we nonetheless thought it worthwhile to test a few variations of the Creative Adversarial Network, as described in Elgammal et al. (2017). Lacking an official repo or weights, we had to base our implementation off of what was described in the paper. The diffusion models disc-full and disc-med used the classification heads of discriminators trained in this fashion for their reward model, but the images being evaluated in this section were created by the generators used in the Creative Adversarial Network. In the original CAN paper, all images were generated at resolution 256, trained on the Full wikiart dataset, with batch size 128 without gradient penalty. However, we wanted to slightly expand our search space, so we tried every possible combination of four choices
  1. 1.
    Image Dimensionality: 256 or 512
  2. 2.
    Dataset: Full or Mediums
  3. 3.
    Batch Size: 256 or 128
  4. 4.
    Gradient Penalty: GP (We do use GP) or No GP (We do not use GP)
This gave us sixteen combinations, and so we had 16 generators. Scores are shown in table 3. Refer to appendix A for box plots visualizing these results.
Model AVA IR
Image Dim 256; Full Dataset
Batch 256 No GP 3.53 ( 0.13 ) -2.23 ( 0.02 )
Batch 256 GP 3.73 ( 0.17 ) -2.14 ( 0.03 )
Batch 128 No GP 3.9 ( 0.16 ) -1.39 ( 0.23 )
Batch 128 GP 3.79 ( 0.16 ) -1.29 ( 0.23 )
Image Dim 512; Full Dataset
Batch 256 No GP 4.1 ( 0.16 ) -2.13 ( 0.05 )
Batch 256 GP 3.64 ( 0.08 ) -2.2 ( 0.02 )
Batch 128 No GP 4.25 ( 0.08 ) -1.78 ( 0.1 )
Batch 128 GP 4.21 ( 0.14 ) -0.89 ( 0.1 )
Image Dim 256; Mediums Dataset
Batch 256 No GP 3.88 ( 0.19 ) -1.46 ( 0.18 )
Batch 256 GP 3.85 ( 0.13 ) -1.6 ( 0.18 )
Batch 128 No GP 4.38 ( 0.23 ) -1.64 ( 0.18 )
Batch 128 GP 4.04 ( 0.14 ) -1.39 ( 0.09 )
Image Dim 512; Mediums Dataset
Batch 256 No GP 4.31 ( 0.24 ) -1.53 ( 0.12 )
Batch 256 GP 3.97 ( 0.14 ) -1.31 ( 0.12 )
Batch 128 No GP 4.34 ( 0.2 ) -1.22 ( 0.13 )
Batch 128 GP 4.23 ( 0.12 ) -1.2 ( 0.14 )
Table 3: Caption
The GANs did not perform well. The highest AVA score attained by a GAN was 4.34, which was lower than the AVA score for all of the diffusion models. The highest IR score was -0.89, which was higher than the IR score for disc-med, clip-full and disc-full, but the average GAN score was still far below that of the diffusion models. Evidently, GANs do not seem to perform very well. Their inferiority is compounded by the problems of training GANs, such as mode collapse and instability (Arjovsky & Bottou, 2017). The much lower standard deviation of aesthetic scores for the GANs also strongly indicates mode collapse.
5.2 User Study
We also conducted a user study to evaluate a subset of the models. The goal of this study was to evaluate each model for human appreciation, as we had done with IR and AVA, as well as novelty, surprise, ambiguity and complexity. For each model we did evaluate in this user study, we generated five images with the same seeds. Then, for each image, users to gauge 5 metrics, by asking them each a question, as was done in Elgammal et al. (2017). These 5 metrics were based on the components of aesthetic arousal (Berlyne, 1971), that being the level of alertness or excitation that an image provokes in a viewer. "Too much" arousal can actually have a negative effect on subjective appreciation. The metrics and their associated questions are as follows:
  1. 1.
    Appreciation: How much do you like this image: 1-extremely dislike, 2-dislike, 3-Neutral, 4-like, 5-extremely like.
  2. 2.
    Novelty: Rate the novelty of the image: 1-extremely not novel, 2-some how not novel, 3-neutral, 4 somehow novel, 5-extremely novel.
  3. 3.
    Surprise: Do you find the image surprising: 1-extremely not surprising, 2-some how not surprising, 3-neutral, 4-some how surprising, 5-extremely surprising.
  4. 4.
    Ambiguity: Rate the ambiguity of the image. I find this image: 1-extremely not ambiguous, 2-some how not ambiguous, 3-neutral, 4-some how ambiguous, 5-extremely ambiguous.
  5. 5.
    Complexity: Rate the complexity of the image. I find this image: 1-extremely simple, 2-some how simple, 3-neutral, 4-somehow complex, 5-extremely complex
We used prolific.com to find users, who were then redirected to a google forms survey. Each user was paid 10.00 dollars. Median time to complete the survey was 12 minutes and 28 seconds.
Appreciation Novelty Surprise Ambiguity Complexity
disc-full 3.048 2.784 2.576 2.920 3.040
clip-full 3.376 2.920 2.728 2.656 3.224
kmeans-full 3.104 2.760 2.656 2.808 3.248
disc-med 2.600 2.728 2.520 3.144 3.064
clip-med 2.936 3.032 2.808 2.784 3.104
kmeans-med 3.216 3.328 2.960 3.448 3.728
Table 4: User Study Results
Results are shown in table 4. We see here that kmeans-med outperforms all of the other models on all of the metrics. For both datasets, the lowest Appreciation score was attained by the baseline discriminator model. For the novelty metric, which is a proxy for the "creative" aspect of creativity, both discriminator-free methods outperform the discriminator using the mediums datasets. For the other metrics, which are less important but still interesting given that they are components of aesthetic arousal, the discriminator methods still underperform. Complexity for example, shows that for both datasets, the discriminator performs the worst.
5.3 Visual Results
We also provide a few visual results in figure 1. Each image in each row was generated with the same prompt and random seed. The leftmost column has the prompt used to generate each picture in each row, and each column is labeled with the model used to generate the images in said column. For more pictures, consult appendix B.
Figure 1: Image Comparisons
Both discriminator model, as well as clip-full generate blurry, abstract patterns. While those are certainly interesting, they do not match the prompt. Kmeans-full also suffers from a proclivity to draw mens’ faces, regardless of prompt. However, both clip-med and kmeans-med demonstrate prompt alignment and interesting visuals.
5.4 Content and Style Similarities
We also compared the similarities between each model. Given that the nth image generated by each model used the same prompt and seed across all models, we could compare the cosine distance of the embeddings of the nth image for each model with the embeddings of the nth image of every other model, and average them. We used style and content embeddings from the dino-vits16 checkpoint (Caron et al., 2021) from https://huggingface.co/facebook/dino-vits16. Many other works have commented on the ability to extract separable style and content from vision transformers (Tumanyan et al., 2022; Kwon & Ye, 2023).
disc-full clip-full kmeans-full disc-med clip-med kmeans-med
disc-full 1.0 0.29 0.26 0.29 0.27 0.29
clip-full 0.29 1.0 0.24 0.41 0.27 0.26
kmeans-full 0.26 0.24 1.0 0.24 0.38 0.31
disc-med 0.29 0.41 0.24 1.0 0.25 0.26
clip-med 0.27 0.27 0.38 0.25 1.0 0.33
kmeans-med 0.29 0.26 0.31 0.26 0.33 1.0
Table 5: Average Content Similarities
disc-full clip-full kmeans-full disc-med clip-med kmeans-med
disc-full 1.0 0.23 0.23 0.22 0.27 0.27
clip-full 0.23 1.0 0.28 0.45 0.32 0.3
kmeans-full 0.23 0.28 1.0 0.21 0.59 0.4
disc-med 0.22 0.45 0.21 1.0 0.26 0.25
clip-med 0.27 0.32 0.59 0.26 1.0 0.49
kmeans-med 0.27 0.3 0.4 0.25 0.49 1.0
Table 6: Average Style Similarities
5.5 Visualizing The Possibility Space
For each model, we combined each pair of evaluation datasets and then then performed dimensionality reduction using PCA and TSNE to embed images into 2 dimensions. Thus, we could visualize the overlap, or lack thereof, between the possibility space of each model.
Figure 2: Possibility Space
This visualization, as well as the similarity tables, proves that all the models being studied are creating images in very distinct spaces.
6 Conclusion
In this paper, we introduce versatile CLIP-based and K-Means-based creative style ambiguity losses, which eliminate the need for training a separate style classifier on labeled datasets. We have demonstrated that these new style ambiguity losses are just as good, oftentimes better, than the baseline discriminator based style ambiguity loss, as shown by automated metrics and user studies. Furthermore, this method can likely be applied to different datasets, consisting of different images or even different modalities like audio or video. Using the K-Means Classifier for style ambiguity loss is particularly adept at this, given there is no need for any labels. However, K-Means is just one unsupervised clustering method, and others may be worth exploring as well.
Broader Impact Statement
Many are concerned about the impacts of generative AI. By making art, this work infringes upon a domain once exclusive to humans. Companies have faced scrutiny for possibly using AI (Gutierrez, 2024), and many creatives, such as screenwriters and actors, have voiced concerns about whether their jobs are safe (del Barco, 2023). Nonetheless, using AI can help humans by making them more efficient, providing inspiration, and generating ideas (Fortino, 2023; Campitiello, 2023; Darling, 2022). It’s also not certain how copyright protection will function for AI-generated art (Watiktinnakorn et al., 2023), given copyright law is based on the premise that creative works originate solely from human authorship. Clear, consistent policies, both at the government level and by industry and/or academic groups, will be needed to mitigate the harm and maximize the benefits for all members of society.
7 Assistance
Author Contributions
This work was done without any outside assistance or collaboration.
Acknowledgements
Rutgers Office of Advanced Research Computing kindly provided the computing infrastructure to run the experiments. A special thanks goes out to Dr. Ahmed Elgammal and Dr. Eugene White for their past advice prior to this work.
Appendix A GAN Results
Box plots for the image reward and aesthetic scores are shown in tables 7 and 8, respectively. Some example images are shown in figure 3
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
Table 7: Image Reward
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
Table 8: Aesthetic Reward
Figure 3: GAN Examples
Appendix B Additional Images
We display a few more images in figures 4 and 5. Once again, each image in each row was generated with the same seed and same prompt.
Figure 4: More Images
Figure 5: More Images
Appendix C WikiArt Classes
We host the original WikiArt-Full dataset on REDACTEDWHILEUNDERBLINDREVIEW. The 27 WikiArt style classes, $L_{full}$ are listed in table 9
contemporary-realism art-nouveau-modern abstract-expressionism
northern-renaissance mannerism-late-renaissance early-renaissance
realism action-painting color-field-painting
pop-art new-realism pointillism
expressionism analytical-cubism symbolism
fauvism minimalism cubism
romanticism ukiyo-e high-renaissance
synthetic-cubism baroque post-impressionism
impressionism rococo na-ve-art-primitivism
Table 9: Styles
Appendix D Medium Subset
D.1 Captions
All captions were generated using the Salesforce/blip-image-captioning-base checkpoint downloaded from https://huggingface.co/Salesforce/blip-image-captioning-base.
D.2 Subset
The categories of images that had the highest percentage of images in the categories with text captions that contained one of the words in the relevant text prompt set, as well as the percentage of images that did so, and the quantity of images in said category are shown in table 10 . We host the Mediums datasets at REDACTEDWHILEUNDERBLINDREVIEW.
style class quantity percentage
expressionism 6054 91.56
post-impressionism 5832 89.25
fauvism 841 96.08
abstract-expressionism 2518 89.95
na-ve-art-primitivism 2148 93.39
cubism 2027 91.61
synthetic-cubism 197 89.85
analytical-cubism 105 91.43
new-realism 280 96.07
action-painting 93 93.55
Table 10: Mediums
Appendix E Training
For reproducibility and transparency, the hyperparameters are listed in table 11 and table 12. All experiments were implemented in Python, building the models in pytorch (Paszke et al., 2017) using accelerate (Gugger et al., 2022) for efficient training. The diffusion models also relied on the trl (von Werra et al., 2020), diffusers (von Platen et al., 2022) and peft (Mangrulkar et al., 2022) libraries. The K-Means clustering was done using the k means implementation from scikit-learn (Pedregosa et al., 2011). A repository containing all code can be found on github at REDACTEDWHILEUNDERREVIEW. Each experiment was run using two NVIDIA A100 GPUs with 40 GB RAM.
Hyperparameter Value
Epochs 25
Effective Batch Size 8
Batches per Epoch 32
Inference Steps per Image 30
LORA Matrix Rank 4
LORA $\alpha$ 4
Optimizer AdamW
Learning Rate 0.0015
AdamW $\beta_{1}$ 0.9
AdamW $\beta_{2}$ 0.99
AdamW Weight decay 1e-4
AdamW $\epsilon$ 1e-8
Table 11: DDPO Hyperparameters
Hyperparameter Value
Epochs 100
Batch Size 32
Optimizer Adam
Learning Rate 0.001
Adam $\beta_{1}$ 0.9
Adam $\beta_{2}$ 0.99
Adam Weight decay 0.0
Adam $\epsilon$ 1e-8
Noise Dim 100
Wasserstein $\lambda$ 10
Leaky ReLU negative slope 0.2
Convolutional Kernel 4
Convolutional Stride 2
Transpose Convolutional Kernel 4
Transpose Convolutional Stride 2
Table 12: CAN Hyperparameters
E.1 Architecture
For diffusion model training, the text encoder, autoencoder and unet were all loaded from https://huggingface.co/stabilityai/stable-diffusion-2-base. These model components were all frozen, but we added trainable LoRA weights to the cross-attention layers of the Unet. Parameter counts are shown in table 13. The diffusion model components used the same amount of parameters regardless of image size, but the generator and discriminator had more parameters as image size increased.
Model Component Total Parameters Trainable Parameters Percent Trainable
Text Encoder 34,0387,840 0 0%
Autoencoder 83,653,863 0 0%
UNet 866,740,676 829,952 0.1%
Generator (Image Dim 512) 48,014,784 48,014,784 100%
Discriminator (Image Dim 512) 20,115,932 20,115,932 100%
Table 13: Parameter Counts
We used the convolutional neural network (Dumoulin & Visin, 2018) architecture described in Elgammal et al. (2017) for the CAN but had to use more/less layers to produce higher/lower dimension images. The generator takes a $1\times 100$ gaussian noise vector $\in\mathbb{R}^{100}\sim\mathcal{N}(0,I)$ and maps it to a $4\times 4\times 2048$ latent space, via a convolutional transpose layer with kernel size = 4 and stride =1, followed by 6 transpose convolutional layers each upscaling the height and width dimensions by two, and halving the channel dimension (for example one of these transpose convolutional layers would map $\mathbb{R}^{4\times 4\times 2048}\rightarrow\mathbb{R}^{8\times 8\times 1024}$) followed by batch normalization (Ioffe & Szegedy, 2015) and Leaky ReLU (Maas et al., 2013), and then one final convolutional transpose layer with output channels = 3 and tanh (Dubey et al., 2022) activation function. Diagrams of the generator is shown in the figure 6.
For the discriminator, we first applied a convolution layer to downscale the input image height width dimensions by 2 and mapped the 3 input channel dimensions to 32 ($\mathbb{R}^{512\times 512\times 3}\rightarrow\mathbb{R}^{256\times 256\times 32}$) with Leaky ReLU activation. Then we had 5 convolutional layers each downscaling the height and width dimensions by 2 and doubling the channel dimension (for example, one of these convolutional layers would map $\mathbb{R}^{256\times 256\times 32}\rightarrow\mathbb{R}^{128\times 128\times 64}$) with batch normalization and Leaky ReLU activation. Then we had two more convolutional layers, each downscaling the height and width dimensions but keeping the channel dimensions constant (using the prior layer’s channel dimensions), with batch normalization and Leaky ReLU activation. The output of the convolutional layers was then flattened. The discriminator had two heads- one for style classification (determining which style a real image belongs to) and one for binary classification (determining whether an image was real or fake). The binary classification head consisted of one linear layer with one output neuron. The style classification layer consisted of 2 linear layers with LeakyReLU activation and Dropout, with output 1024 output neurons and 512 output neurons, respectively, followed by a linear layer with 27 output neurons for the 27 artistic style classes. A diagram of the discriminator is shown in 7.
Figure 6: Generator Architecture (Image Dim 512)
Figure 7: Discriminator Architecture (Image Dim 512)