### v3

# Style Ambiguity Loss Using CLIP

Abstract
In this work, we explore using the style ambiguity training objective, originally used to approximate creativity, on a diffusion model. However, this objective requires the use of a pretrained classifier and a labeled dataset. We introduce new forms of style ambiguity loss that do not require training a new classifier or a labeled dataset. Instead of using a classifier, we generate centroids in the CLIP embedding space, and images are classified based on their relative distance to said centroids. We find the centroids via K-means clustering of an unlabeled dataset, as well as using text labels to generate CLIP embeddings, to be used as centroids. Code is available at https://github.com/jamesBaker361/clipcreate
Index Terms:
CLIP, multimodal, diffusion
I Introduction
Much of generative imagery is trained to mimic the underlying dataset. Generating truly ”creative” work was initiated by the Creative Adversarial Network [1], which used a style ambiguity loss to train a generator network to generate images that could not be classified as belonging to a particular style. However, every set of styles or concepts requires training a classifier before even training a model to generate images, which itself takes time and requires a labeled dataset, the collection and curation of which can be expensive. To circumvent these issues, we propose using a classification technique for style ambiguity that does not require any additional training and can be easily applied to any dataset, labeled or unlabeled. We demonstrate this by using reinforcement learning to finetune diffusion models with the various style ambiguity losses. Our contributions are as follows:
  • We developed versatile Text-based and K-Means-based creative style ambiguity losses that do not require training a separate style classifier on a labeled dataset, instead using CLIP for generating centroids.
  • We find that diffusion models trained with our novel forms of style ambiguity loss score similar or higher on automated metrics of human appreciation to diffusion models trained without, and perform a user study to show that our novel models also score higher on components of creativity.
II Background
II-A Creative Adversarial Network
A Generative Adversarial Network, or GAN [2], consists of two models, a generator and a discriminator. The generator generates samples from noise, and the discriminator detects if the samples are drawn from the real data or generated. During training, the generator is trained to trick the discriminator into classifying generated images as real, and the discriminator is trained to classify images correctly. Given a generator $G:\mathbb{R}^{noise}\mathbb{R}ightarrow\mathbb{R}^{h\times w\times 3}$, a discriminator $D:\mathbb{R}^{h\times w\times 3}\mathbb{R}ightarrow[0,1]$ real images $x\in\mathbb{R}^{h\times w\times 3}$, and noise $\mathcal{Z}\in\mathbb{R}^{noise}$, the objective is:
$\underset{G}{min}\>\underset{D}{max}\>\mathbb{E}_{x}[log(D(x)]+\mathbb{E}_{\mathcal{Z}}[log(1-D(G(\mathcal{Z}))]$
[1] introduced the Creative Adversarial Network, or CAN, which was a DCGAN [3] where the discriminator was also trained to classify real samples, minimizing the style classification loss. Given $N$ classes of image (such as ukiyo-e, baroque, impressionism, etc.), the classification modules of the Discriminator $D_{C}:\mathbb{R}^{h\times w\times 3}\mathbb{R}ightarrow\mathbb{R}^{N}$ that returns a probability distribution over the $N_{s}$ style classes for an image and the real labels $\ell\in\mathbb{R}^{N}$, the style classification loss was:
$L_{SL}=\mathbb{E}_{x,\ell}[\textbf{CE}(D_{C}(x),\ell)]$
Where CE is the cross entropy function.
The generator was also trained to generate samples that could not be easily classified as belonging to one class. This stylistic ambiguity is a proxy for creativity or novelty. Given a vector $U\in\mathbb{R}^{N}$, where each entry $u_{1},u_{2},,,u_{N}=\frac{1}{N}$, and the classification modules of the discriminator $D_{C}$ the style ambiguity loss is:
$L_{SA}=\mathbb{E}_{\mathcal{Z}}[\textbf{CE}(C(G(\mathcal{Z})),U)]$
The discriminator was additionally trained to minimize $L_{SL}$ and the generator was additionally trained to minimize $L_{SA}$.
II-B Diffusion
A diffusion model aims to learn to iteratively remove the noise from a corrupted sample to restore the original. Starting with $x_{0}$, the forward process $q$ iteratively adds Gaussian noise to produce the noised version $x_{T}$, using a noise schedule $\beta_{1}...\beta_{T}$, which can be learned or manually set as a hyperparameter:
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I})$
More importantly, we also want to model the reverse process $p$, that turns a noisy sample $x_{T}$ back into $x_{0}$, conditioned on some context $c$. As $x_{T}$ is the fully noised version, $p(x_{T}|c)=\mathcal{N}(x_{T};\mathbf{0},\mathbf{I})$
$p_{\theta}(x_{t-1}|x_{t},c)=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t,c),\Sigma_{\theta}(x_{t},t,c))$
Once the model has been trained, the reverse process, aka inference, generates a sample from noise $x_{T}\sim\mathcal{N}(0,1)$. We use the DDIM technique [4] for sampling. In this work, we the more popular variant, Stable Diffusion [5], where $x$ is replaced with a latent embedding $\mathcal{E}(x)$, where $\mathcal{E}:\mathbb{R}^{h\times w\times 3}\mathbb{R}ightarrow\mathbb{R}^{h_{z}\times w_{z}\times c_{z}}$ is a frozen autoencoder [6], and $h_{z}<h,w_{z}<w,c_{z}>3$.
III Methods
III-A Denoising Diffusion Proximal Optimisation
Introduced by [7], Denoising Diffusion Proximal Optimisation, or DDPO, represents the reverse Diffusion Process as a Markov Decision Process [8]. A similar method was also pursued by [9]. Reinforcement learning training was then applied to a pretrained diffusion model. Following [10], [7] also implemented clipping to protect the policy gradient from excessively large updates, and per prompt stat tracking to normalize rewards. We largely follow their method but use a different reward function. We fine-tune off of the pre-existing stabilityai/stable-diffusion-2-base checkpoint [5] downloaded from https://huggingface.co/stabilityai/stable-diffusion-2-base.
III-B Text Prompts
DDPO does not require any new data, given that we are fine-tuning off of a pretrained checkpoint. However, each time we train the model, we must decide which text prompts to use to condition the generation of images. We used the set of (painting, drawing, art) as our text prompts.
III-C Datasets and Labels
Training the CAN, of course, requires a labeled dataset. We used two different real datasets based off of WikiArt [11]:
  1. 1.
    Full: the WikiArt dataset as is. Consists of roughly 80k images.
  2. 2.
    Mediums: We used BLIP [12] to generate captions, and then selected the 10 classes that had the highest fraction of their descriptions containing any of the words in the text prompt set (painting, drawing, art). We then used the images in WikiArt that belonged to those classes. Consists of roughly 20k images.
This also meant we had two different sets of style class labels: $L_{full}$, all 27 class labels, used with Full and $L_{med}$, the 10 labels used in Mediums.
III-D DDPO Reward Function
In the original DDPO paper, the authors used four different reward functions for four different tasks. In this paper, we use the reward model based on stylistic ambiguity, combined with a reward for utility. Given a pretrained CLIP [13] model, that can return a similarity score for each image-text pair: $CLIP:\mathbb{R}^{text}\times\mathbb{R}^{h\times w\times 3}\mathbb{R}ightarrow\mathbb{R}$, image $x_{0}\in\mathbb{R}^{h\times w\times 3}$ generated with text prompt $s$, cross entropy CE, uniform distribution $U\in\mathbb{R}^{N}$ and a classifier $C:\mathbb{R}^{h\times w\times 3}\mathbb{R}ightarrow\mathbb{R}^{N}$ we want to maximize:
$R(x_{0})=\lambda_{novelty}\textbf{CE}(C(x_{0}),U)+\lambda_{utility}CLIP(s,x_{0})$
The first term on the left side of the equation represents style ambiguity loss, and the second term maintains alignment between text and image, essentially keeping the model from straying ”too far” from the text prompt; these terms approximate novelty and utility, respectively. We actually have multiple choices of classifier, which we discuss. The Discriminator Classifier is what was used for the CAN, and the Text-Based Style Classifier and K-Means Image Based Classifer have never been used for style ambiguity loss before, and our contribution is to introduce them and show their merits. Many other works [14, 15] have incorporated CLIP embeddings as a condition to generate images, but we only use the CLIP classifier to generate a reward.
III-D1 Discriminator Classifier
We can use the classification module of the CAN discriminator as the classifier in the reward function, setting $C=D_{C}$. This is the baseline against which we are comparing the other two types of classifier with. This discriminator was trained at image resolution 512, with batch size 128. Given that we had two datasets (Full and Mediums), we actually had to train two of these discriminators, one trained for each.
III-D2 Text-Based Style Classifier
For each generated image $x_{0}$, for each class name $s_{i},1\leq i\leq N_{s}$ in the style class label set we want to use, we find $CLIP(s_{i},x_{0})$. We can then create a vector $(CLIP(s_{1},x_{0}),CLIP(s_{2},x_{0}),,,CLIP(s_{N_{s}},x_{0}))$ and then use softmax to normalize the vector and define the result as $TextC(x_{0})$. Formally:
$TextC(x_{0})=\textbf{softmax}((CLIP(s_{1},x_{0}),,,CLIP(s_{N_{s}},x_{0}))$
Then we set $C=TextC$. Given the two sets of labels, $L_{full}$ and $L_{med}$, we had two different types of Text-Based classifier.
III-D3 K-Means Image Based Classifiers
[…]
Then we set $C=KM$. Given two datasets (Full and Mediums), we had two sets of centers.
IV Results
Given three types of classifier and two datasets, we had six models. Each model was trained for 100 epochs on a single NVIDIA A100 Tensor Core GPU, elapsing roughly 14 hours each. Note that for methods text-full and text-med, we don’t use the data in the data for the classifier per se; however, we use the style class labels, which are unique to the datasets in question. text-full uses $L_{full}$ and text-med uses $L_{mediums}$. We generated all images with width and height = 512.
IV-A Automated Evaluation
For each dataset, for each choice of classifier, we made an evaluation dataset of 200 images. We used the exact same prompts and random seeds for each. I.e. if the nth image generated by disc-full used prompt ”painting” and random seed = $z\in\mathbb{Z}$, so would the nth image generated by text-full, kmeans-full, etc. for prompts, we had 24 prompts, each generated by concatenating one of [”painting”,”art”,”drawing”] with one of [” of a man ”, ” of a woman ”,” of nature ”,” of an animal ”, ” of a city”,” ”,” of a building ”,” of a child ”] We used two scoring metrics to score the models. Both of these metrics are metrics of human appreciation:
  • AVA Score: (AVA) The AVA model was trained on the AVA dataset [16] of images and average rankings by human subjects, in order to learn to approximate human preferences given an image. We used the CLIP model weights from the clip-vit-large-patch14 checkpoint and the Multi-Layer Perceptron weights downloaded from https://huggingface.co/trl-lib/ddpo-aesthetic-predictor.
  • Image Reward: (IR) The image reward model [17] was trained to score images given their text description based on a dataset of images and human rankings. We used the image-reward python library found at https://github.com/THUDM/ImageReward/tree/main.
Model AVA IR
discfulldisc-full 4.96 (0.5) -1.26 (1.04)
clipfullclip-full 4.99 (0.46) -1.59 (0.63)
kmeansfullkmeans-full 5.71 (0.42) 0.03 (1.35)
discmeddisc-med 5.12 (0.4) -1.55 (0.59)
clipmedclip-med 5.44 (0.41) 0.91 (0.89)
kmeansmedkmeans-med 5.09 (0.67) 0.77 (0.87)
TABLE I: Aesthetic Scores, Diffusion Models
Results of our experiments are shown in Table I Each cell contains the mean (and standard deviation). For both datasets, for both metrics, either a text model or kmeans model scored higher than the baseline discriminator model. This shows that our discriminator-free approaches in do not come at the cost of lower quality, and in fact are likely to produce better images than the discriminator baseline.
IV-B User Study
We also conducted a user study to evaluate a subset of the models. The goal of this study was to evaluate each model for human appreciation, as we had done with IR and AVA, as well as novelty, surprise, ambiguity and complexity. For each model we did evaluate in this user study, we generated five images with the same seeds. Then, for each image, users to gauge 5 metrics, by asking them each a question, as was done in [1]. These 5 metrics were based on the components of aesthetic arousal [18], that being the level of alertness or excitation that an image provokes in a viewer. ”Too much” arousal can actually have a negative effect on subjective appreciation. The metrics and their associated questions are as follows:
  1. 1.
    Appreciation (App): How much do you like this image: 1-extremely dislike, 2-dislike, 3-Neutral, 4-like, 5-extremely like.
  2. 2.
    Novelty (Nov): Rate the novelty of the image: 1-extremely not novel, 2-some how not novel, 3-neutral, 4 somehow novel, 5-extremely novel.
  3. 3.
    Surprise (Sur): Do you find the image surprising: 1-extremely not surprising, 2-some how not surprising, 3-neutral, 4-some how surprising, 5-extremely surprising.
  4. 4.
    Ambiguity (Amb): Rate the ambiguity of the image. I find this image: 1-extremely not ambiguous, 2-some how not ambiguous, 3-neutral, 4-some how ambiguous, 5-extremely ambiguous.
  5. 5.
    Complexity (Com): Rate the complexity of the image. I find this image: 1-extremely simple, 2-some how simple, 3-neutral, 4-somehow complex, 5-extremely complex
We used prolific.com to find users, who were then redirected to a google forms survey. Each user was paid 10.00 dollars. Median time to complete the survey was 12 minutes and 28 seconds.
App Nov Sur Amb Com
disc-full 3.048 2.784 2.576 2.920 3.040
clip-full 3.376 2.920 2.728 2.656 3.224
kmeans-full 3.104 2.760 2.656 2.808 3.248
disc-med 2.600 2.728 2.520 3.144 3.064
clip-med 2.936 3.032 2.808 2.784 3.104
kmeans-med 3.216 3.328 2.960 3.448 3.728
TABLE II: User Study Results
Results are shown in table II. We see here that kmeans-med outperforms all of the other models on all of the metrics. For both datasets, the lowest Appreciation score was attained by the baseline discriminator model. For the novelty metric, which is a proxy for the ”creative” aspect of creativity, both discriminator-free methods outperform the discriminator using the mediums datasets.
IV-C Visual Results
We also provide a few visual results in figure 1. Each image in each row was generated with the same prompt and random seed. The leftmost column has the prompt used to generate each picture in each row, and each column is labeled with the model used to generate the images in said column.
Both discriminator model, as well as text-full generate blurry, abstract patterns. While those are certainly interesting, they do not match the prompt. However, both text-med and kmeans-med demonstrate prompt alignment and interesting visuals.
Figure 1: Image Comparisons
V Conclusion
In this paper, we introduce versatile CLIP-based and K-Means-based creative style ambiguity losses, which eliminate the need for training a separate style classifier on labeled datasets. We have demonstrated that these new style ambiguity losses are just as good, oftentimes better, than the baseline discriminator based style ambiguity loss, as shown by automated metrics and user studies. This method can be applied to different image datasets or even different modalities like audio or video. Using the K-Means Classifier, or any unsupervised clustering method, for style ambiguity loss is particularly adept at this, given there is no need for any labels.